# Kac–Moody Symmetric Spaces: arbitrary symmetrizable complex or almost split real type

### v4

@@ 2 Basics on Symmetric Spaces
[…]
@@
[…]
@@
$\left(\alpha\circ s_{x}\circ\alpha^{-1}\right)(p)=\alpha\left(s_{x}\left(\alpha^{-1}(p)\right)\right)=\alpha\left(\mu\left(x,\alpha^{-1}(p)\right)\right)=\mu(\alpha(x),p)=s_{\alpha(x)}(p).$
@@
[…]
@@ 3.1 Kac–Moody Algebra
The original approach of Kac is to define that the sets $\Pi$ and $\Pi^{\vee}$ are linearly independent, see [Kac90, §1.1]. In the general case where the generalized Cartan matrix $\mathbb{A}$ is not necessarily invertible (e.g. if $\mathbb{A}$ is of affine type), we have that $\Pi$ is not linearly independent. To address this problem, one can extend the Cartan subalgebra $\mathfrak{h}$ to ensure that the $\alpha_{i}$ are linearly independent. For a concrete discussion of this, see [Kac90, Proof of Proposition 1.1] and [Mar18, Chapter 3.5 and Example 7.10]. It turns out that for $\mathbb{A}\in\mathbb{Z}^{n\times n}$, where $l$ denotes the rank of $\mathbb{A}$, the extended Cartan subalgebra must have dimension $2n-l$ and is isomorphic to $\mathbb{C}^{2n-l}$. The dual space is also extended and constructed in such a way that the simple roots still have the relation $\langle\alpha_{i},\alpha_{j}^{\vee}\rangle=a_{ij}\in\mathbb{A}$.
On this basis, we denote the following as $\mathcal{D}_{\mathrm{Kac}}^{\mathbb{A}}$ the root datum where the elements $c_{i}$ and $h_{i}$ are linearly independent and denote by $\Lambda_{\mathrm{Kac}}$ the corresponding $\mathbb{Z}$-module, which has minimum rank with respect to these properties, i.e. the rank of $\Lambda_{\mathrm{Kac}}$ is equal to $2n-l$. Similarly, we use $\Lambda_{\mathrm{Kac}}^{\vee}$ to denote the $\mathbb{Z}$-dual. Moreover, denote by $\{v_{i}\mid 1\leq i\leq 2n-l\}$ the $\mathbb{Z}$-basis of the extended $\Lambda_{\mathrm{Kac}}$, where $h_{i}=v_{i}$ for $1\leq i\leq n$. We follow here the notation and idea of [Mar18, Example 7.10].
@@
Denote by $\Delta$ all roots of $\mathfrak{g}(\mathbb{A})$. Furthermore, a root $\alpha$ is called positive if the coefficients this representation are non-negative, and negative otherwise. By [Mar18, Proposition 3.14 (5)] follows that every root is either positive or negative.
@@ 3.2 The Tits Functor
Since $\{v_{i}\}$ is a $\mathbb{Z}$-basis of $\Lambda_{\mathrm{Kac}}^{\vee}$ (see 3.3) and the rank is equal $2n-l$, the set
[…]
@@
[…]
@@
$U_{\alpha_{i}}\coloneqq\varphi_{i}\left(\left\{\begin{pmatrix}1&r\\ 0&1\end{pmatrix}\middle|\;r\in\mathbb{K}\right\}\right)\;\text{ and }\;\;U_{-\alpha_{i}}\coloneqq\varphi_{i}\left(\left\{\begin{pmatrix}1&0\\ -r&1\end{pmatrix}\middle|\;r\in\mathbb{K}\right\}\right).$
@@
$\tilde{s}_{i}\coloneqq\varphi_{i}\begin{pmatrix}0&1\\ -1&0\end{pmatrix}.$
@@
$x_{\alpha_{i}}(r)\coloneqq\varphi_{i}\begin{pmatrix}1&r\\ 0&1\end{pmatrix}\in U_{\alpha_{i}},$
@@
$x_{-\alpha_{i}}(r)\coloneqq\varphi_{i}\begin{pmatrix}1&0\\ -r&1\end{pmatrix}\in U_{-\alpha_{i}},$
@@
Now let $\mathbb{B}\in\mathbb{Z}^{2n-l\times 2n-l}$ be a generalized Cartan matrix with rank $2n-l$, such that $\mathbb{A}$ is included in $\mathbb{B}$ as a principal submatrix. This can be done in a similar way to expanding a non-invertible generalized Cartan matrix to obtain the extended Cartan algebra, see [Kac90, Proposition 1.1] or [Mar18, Section 3.5] (or 3.3). Consider now the corresponding root datum
@@ 3.3 Twin Buildings
The pair $(\mathcal{W},S)$ with generating set $S=\{r_{1},...,r_{n}\}$ is a Coxeter system. Note that any Weyl group of a Kac–Moody algebra forms a Coxeter system and the order of the elements are in relation with the entries of the generalized Cartan matrix, see [Mar18, Proposition 4.22]. In detail, any finite Coxeter group is called a Weyl group.
@@
Now let $G$ be a group with subgroup $B$, $(\mathcal{W},S)$ a Coxeter system and $C\colon\mathcal{W}\to B\backslash G/B$ the bijection given by
@@
Examples of groups that possess a (twin) $BN$-pair are the general linear group ([AB08, Chapter 6.5]), the symplectic group ([AB08, Chapter 6.6]), orthogonal groups ([AB08, Chapter 6.7]) and unitary groups ([AB08, Chapter 6.8]).
@@
    For all $\alpha,\beta\in\Delta^{\mathrm{re}}$ with $\alpha\neq\beta$ such that the set $\{\alpha,\beta\}$ is prenilpotent, we have $[U_{\alpha},U_{\beta}]\leq U_{(\alpha,\beta)}$, where $(\alpha,\beta)\coloneqq[\alpha,\beta]\setminus\{\alpha,\beta\}$. For details on the notation see [AB08, Definition 8.41].
@@ 3.4 A Characteristic Subgroup
A very useful subgroup for studying symmetric spaces associated to Kac–Moody group is the centered subgroup, that is the subgroup only generated by the root groups.
@@
Since the RGD axioms are manly using root groups, we can deduce that $(G^{\dagger},(U_{\alpha})_{\alpha\in\Delta^{\mathrm{re}}},T^{\dagger})$ forms an RGD system. Moreover, according to [AB08, Proposition 8.82], the kernel of the action of $G$ on the associated twin building $\Delta_{\pm}$ is contained in the torus $T$, so that
@@
cf. [AB08, Corollary 8.79]. Since the natural map $T\mapsto G/G^{\dagger}$ is surjective, the quotient $G/G^{\dagger}$ is abelian and thus $[G,G]\subseteq G^{\dagger}$.
@@
$[G,G]\subseteq G^{\dagger}=[G^{\dagger},G^{\dagger}]\subseteq[G,G]$
@@ 3.5 The Chevalley Involution
The statement of (A2) follows directly from the definition of the maps, see the paragraph above. Concretely, this shows that $A=\tau(T)\cong(\mathbb{R}_{>0})^{2n-l}$, because the dimension of $T$ is $2n-l$.
@@
Finally, we prove the statement of (A5). Therefore recall, that by the definition of $B_{\pm}$ one has $\Theta(B_{\pm})=B_{\mp}$ and that $T=B_{+}\cap B_{-}$. Note, that this is based on the fact that $U_{+}\cap U_{-}=\{e\}$. Thus we have $K\cap B_{\pm}=K\cap T$ = M, by definition. ∎∎
@@ 3.6 The Iwasawa Decomposition
since $M,K^{\dagger}\subset K$ (see 3.14 (A5)). The other order of the decomposition results from the symmetry arguments.
@@
is trivial, i.e., a topological direct product of the fiber and the base space.
@@
Here we recall two lemmas about $k_{\omega}$-spaces for the reader’s convenience that we use in order to derive the topological Iwasawa decomposition.
@@
Lemma 3.18 (Lemma 3.35 from [FHHK20])
Let $(X,(X_{n}))$ be a $k_{\omega}$-pair and let $\pi\colon\widetilde{X}\to X$ be a finite-sheeted covering. Then $(\widetilde{X},\pi^{-1}(X_{n}))$ is a $k_{\omega}$-pair.
@@
is a $k_{\omega}$-pair. Note that the preimage of $D_{l,\pm}$ under the fibration $p_{\pm}$ is exactly $\widetilde{D}_{l,\pm}$ and that the fibration by 3.16 is trivial, so we can apply 3.18, which works here, since the statement is just a topological assertion about $k_{\omega}$-pairs, so we observe that
@@
To conclude that $(K,(K_{l}^{\pm}))$ is a $k_{\omega}$-pair, we must show that the subsets $(K_{l}^{\pm})$ are compact. To this end, recall that $G$ can be written as $KB_{\pm}$ using 3.15, and hence $K$ acts chamber transitively on the to $G$ associated twin building $\Delta_{\pm}$. In particular
@@
This is a finite union of compact sets, see [HKM13, Corollary 3.10]. Recall also that the finite group $M$ is the maximal compact subgroup of $T$ and that $K_{l}^{\pm}\cap B_{\pm}\subseteq M$. Thus
@@
is a finite covering, and so are
@@
$k(g)\coloneqq\iota_{\pm}^{-1}\left(gAU_{\pm}\right)\;\text{ and }\;h(g)\coloneqq k(g)^{-1}g,$
where the maps $\iota_{\pm}$ are the same as in 3.19. Therefore we define the map
@@ 4.1 A criterion for symmetric spaces
Further we call a twin apartment $\Theta$-stable, if the to $G$ associated twin building is invariant as a set under the action of $\Theta$. Recall that $\Delta_{\pm}=G/B_{\pm}$, then we consider the action of $\Theta$ on the building as follows
@@ 4.2 Kac–Moody Symmetric Space
    and it can be deduced that the entries of $D$ are ±$\pm 1$. But by the positive semi-definite property only $+1$ is possible.
@@ 4.3 The Group Model
$\left\langle\tau\left(G^{\dagger}\right)\right\rangle=G^{\dagger}.$
@@
$=\ell_{g}\left(\Theta(g^{-1}h)K\right)$
$=\ell_{g}\left(s_{eK}\left(g^{-1}hK\right)\right)$
$=\left(\ell_{g}\circ s_{eK}\circ\ell_{g^{-1}}\right)(hK).$
@@
$\tau(G)^{2}C_{K}\left(G^{\dagger}A\right).$
@@ 4.4 Flats
$\mathfrak{a}\coloneqq\operatorname{span}_{\mathbb{R}}\left(\alpha^{\vee}_{1},\ldots,\alpha^{\vee}_{n},h_{n+1},\ldots,h_{2n-l}\right)\cong\mathbb{R}^{2n-l}.$
@@
The last step is to check that $gAK\subset G/K$ is closed. 3.20 tells that the multiplication $m\colon U_{\pm}\times A\times K\to G$ is a homeomorphism. This yields that the subspace $AK$ and $gAK$, for any $g\in G$, are closed in $G/K$ since the multiplication with an element induces a homeomorphism within the topological group. ∎
@@ 4.4.1 Flats in the Group Model
Recall that in the case $\mathbb{K}=\mathbb{C}$ $T$ is connected and in the case $\mathbb{K}=\mathbb{R}$ $A$ is the identity component of $T$. This indicates that $N_{K}(T)<N_{K}(A)$, based on the fact that the conjugation is continuous.
@@
To see that this map is well-defined, we first look at $N_{G}(T)$. We know by [FHHK20, Lemma 3.27] that $N_{G}(T)=A\ltimes N_{K}(T)$, since the proof of the quoted statement is given by 3.14 and the refined Birkhoff decomposition (see [Kum02, Theorem 5.2.3 (g)] or proof of 3.21). Together with 4.11 and [FHHK20, Lemma 3.27] follows that $N_{G}(T)$ stabilizes $AK$: Let $g\in N_{G}(T)$, then it can be written as $g=ak$ for $a\in A$ and $k\in N_{K}(A)$ and hence one calculates

### v5

# label=., leftmargin=50pt, rightmargin=10pt label=(), leftmargin=15pt, rightmargin=15pt

@@ 2 Basics on Symmetric Spaces
Since we are interested in generalizations of Riemannian symmetric spaces of non-compact type, the class of symmetric spaces that we consider in this article will automatically satisfy $\textbf{S.4}_{\text{global}}$ instead of just S.4. Note that this in fact allows one to model symmetric spaces on sets instead of topological spaces.
@@
    An example of a Riemannian symmetric space of non-compact type is the hyperbolic space. Since there are no closed geodesics, the property $\textbf{S.4}_{\text{global}}$ is satisfied.
@@
$\left(\alpha\circ s_{x}\circ\alpha^{-1}\right)\hskip-2.0pt(p)=\alpha\hskip-2.0pt\left(s_{x}\hskip-2.0pt\left(\alpha^{-1}(p)\right)\right)=\alpha\hskip-2.0pt\left(\mu\hskip-2.0pt\left(x,\alpha^{-1}(p)\right)\right)=\mu(\alpha(x),p)=s_{\alpha(x)}\hskip-1.0pt(p).$
@@
is surjective. Since $t_{x}\circ s_{b}=s_{x}$ and since $s_{x}=s_{y}$ implies $x=y$ (due to axiom $\textbf{S.4}_{\text{global}}$), the map $\eta$ is also injective. A direct calculation shows that $\eta\colon(X,\mu)\to(T(X,b),\mu_{T})$ is an isomorphism of symmetric spaces. Indeed, one has to check that the reflection map of $(X,\mu)$ is preserved, i.e. $\eta(\mu(x,y))=\mu_{T}(t_{x},t_{y})$. On the right-hand side one can calculate
@@ 3.1 Kac–Moody Algebra
The original approach of Kac is to define that the sets $\Pi$ and $\Pi^{\vee}$ are linearly independent, see [Kac90, §1.1]. In the general case, where the generalized Cartan matrix $\mathbb{A}$ is not necessarily invertible (e.g. if $\mathbb{A}$ is of affine type), the consequence is that $\Pi\subset\mathfrak{h}^{*}$ and $\Pi^{\vee}\subset\mathfrak{h}$ cannot be assumed to be linearly independent.
To address this problem, one can extend the Cartan subalgebra $\mathfrak{h}$, if necessary, to ensure that the $\alpha_{i}$ become linearly independent. For a concrete discussion of this, see [Kac90, Proof of Proposition 1.1] and [Mar18, Chapter 3.5 and Example 7.10]. It turns out that for $\mathbb{A}\in\mathbb{Z}^{n\times n}$, where $l$ denotes the rank of $\mathbb{A}$, the extended Cartan subalgebra must have dimension 2nl2n-l and is isomorphic to $\mathbb{C}^{2n-l}$. The dual space is also extended and constructed in such a way that the simple roots still have the relation $\langle\alpha_{i},\alpha_{j}^{\vee}\rangle=a_{ij}\in\mathbb{A}$.
On this basis, we denote the following as $\mathcal{D}_{\mathrm{Kac}}^{\mathbb{A}}$ the root datum where the elements $c_{i}$ and $h_{i}$ are linearly independent and denote by $\Lambda_{\mathrm{Kac}}$ the corresponding $\mathbb{Z}$-module, which has minimum rank with respect to these properties, i.e. the rank of $\Lambda_{\mathrm{Kac}}$ is equal to 2nl2n-l. Similarly, we use $\Lambda_{\mathrm{Kac}}^{\vee}$ to denote the $\mathbb{Z}$-dual. Moreover, denote by $\{v_{i}\mid 1\leq i\leq 2n-l\}$ the $\mathbb{Z}$-basis of the extended $\Lambda_{\mathrm{Kac}}$, where $h_{i}=v_{i}$ for $1\leq i\leq n$. We follow here the notation and idea of [Mar18, Example 7.10].
@@
Denote by $\Delta$ all roots of $\mathfrak{g}(\mathbb{A})$. Furthermore, a root $\alpha$ is called positive if the coefficients of this representation are non-negative, and negative otherwise. By [Mar18, Proposition 3.14 (5)] follows that every root is either positive or negative.
@@ 3.2 The Tits Functor
Since $\{v_{i}\}$ is a $\mathbb{Z}$-basis of $\Lambda_{\mathrm{Kac}}^{\vee}$ (see 3.3) and the rank is equal 2nl2n-l, the set
$\{r^{v_{i}}\mid r\in R^{\times}\,,\;1\leq i\leq 2n-l\}$
@@
A basis of type $\mathcal{D}_{\mathrm{Kac}}^{\mathbb{A}}$ (as in [Tit87, Section 2], also [Mar18, Definition 7.77]) is defined as a triple $\mathcal{F}\coloneqq\left(\mathcal{G},(\varphi_{i})_{i\in I},\eta\right)$ where $\mathcal{G}\colon\mathbb{Z}\text{-alg}\to\textbf{Grp}$ is a group functor, $(\varphi_{i})_{i\in I}$ is a assortment of morphisms of functors $\varphi_{i}\colon\operatorname{SL}_{2}\to\mathcal{G}$ and a morphism of functors $\eta\colon T\to\mathcal{G}$. One calls the group functor $\mathcal{G}$ Tits functor of type $\mathcal{D}$ if its satisfies the following axioms:
@@
[…]
@@
$\tilde{s}_{i}\coloneqq\varphi_{i}\hskip-2.0pt\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$.
@@
$x_{\alpha_{i}}(r)\coloneqq\varphi_{i}\hskip-2.0pt\begin{pmatrix}1&r\\ 0&1\end{pmatrix}\in U_{\alpha_{i}},$
@@
$x_{-\alpha_{i}}(r)\coloneqq\varphi_{i}\hskip-2.0pt\begin{pmatrix}1&0\\ -r&1\end{pmatrix}\in U_{-\alpha_{i}},$
@@
Now let $\mathbb{B}\in\mathbb{Z}^{2n-l\times 2n-l}$ be a generalized Cartan matrix with rank 2nl2n-l, such that $\mathbb{A}$ is included in $\mathbb{B}$ as a principal submatrix. This can be done in a similar way to expanding a non-invertible generalized Cartan matrix to obtain the extended Cartan algebra, see [Kac90, Proposition 1.1] or [Mar18, Section 3.5] (or 3.3). Consider now the corresponding root datum
@@ 3.3 Twin Buildings
The pair $(\mathcal{W},S)$ with generating set $S=\{r_{1},...,r_{n}\}$ is a Coxeter system.
Note that any Weyl group of a Kac–Moody algebra forms a Coxeter system and the order of the elements are in relation with the entries of the generalized Cartan matrix, see [Mar18, Proposition 4.22]. Not all Coxeter groups appear as Weyl groups, for instance the finite Coxeter group of type $H_{3}$ cannot be a Weyl group.
@@
[…]
@@
Examples of groups that possess a (twin) $BN$-pair are the general linear group ([AB08, Chapter 6.5]), the symplectic group ([AB08, Chapter 6.6]), isotropic orthogonal groups ([AB08, Chapter 6.7]) and isotropic unitary groups ([AB08, Chapter 6.8]).
@@
    For all $\alpha,\beta\in\Delta^{\mathrm{re}}$ with $\alpha\neq\beta$ such that the set $\{\alpha,\beta\}$ is prenilpotent, we have $U_{\alpha},U_{\beta}]\leq U_{(\alpha,\beta)}$, where $(\alpha,\beta)\coloneqq[\alpha,\beta]\setminus\{\alpha,\beta\}$. For details on the notation see [AB08, Definition 8.41].
@@ 3.4 A Characteristic Subgroup
A very useful subgroup for studying symmetric spaces associated to Kac–Moody group is the centered subgroup, that is the subgroup generated only by the root groups.
@@
Since the RGD axioms are mainly using root groups, we can deduce that $(G^{\dagger},(U_{\alpha})_{\alpha\in\Delta^{\mathrm{re}}},T^{\dagger})$ forms an RGD system. Moreover, according to [AB08, Proposition 8.82], the kernel of the action of $G$ on the associated twin building $\Delta_{\pm}$ is contained in the torus $T$, so that
@@
cf. [AB08, Corollary 8.79]. Since the natural map $T\mapsto G/G^{\dagger}$ is surjective, the quotient $G/G^{\dagger}$ is abelian and thus $G,G]\subseteq G^{\dagger}$.
@@
$G,G]\subseteq G^{\dagger}=[G^{\dagger},G^{\dagger}]\subseteq[G,G]$
@@ 3.5 The Chevalley Involution
The statement of (A2) follows directly from the definition of the maps, see the paragraph above. Concretely, this shows that $A=\tau(T)\cong(\mathbb{R}_{>0})^{2n-l}$, because the dimension of $T$ is 2nl2n-l.
@@
Finally, we prove the statement of (A5). Therefore recall, that by the definition of $B_{\pm}$ one has $\Theta(B_{\pm})=B_{\mp}$ and that $T=B_{+}\cap B_{-}$. Note, that this is based on the fact that $U_{+}\cap U_{-}=\{e\}$. Thus we have $K\cap B_{\pm}=K\cap T$ = M, by definition. ∎
@@ 3.6 The Iwasawa Decomposition
since $M,K^{\dagger}\subset K$ (see 3.14 (A5)). The other order of the decomposition holds by symmetry (e.g., via group inversion).
@@
is trivial with compact fiber $M$, i.e., a topological direct product of the fiber and the base space.
@@
Here we recall a lemma about $k_{\omega}$-spaces for the reader’s convenience that we use in order to derive the topological Iwasawa decomposition.
@@
is a $k_{\omega}$-pair. Note that the preimage of $D_{l,\pm}$ under the fibration $p_{\pm}$ is exactly $\widetilde{D}_{l,\pm}$ and that the fibration by 3.16 is trivial with compact fiber $M$, so we observe that
@@
To conclude that $(K,(K_{l}^{\pm}))$ is a $k_{\omega}$-pair, we must show that the subsets $(K_{l}^{\pm})$ are compact. To this end, recall that $G$ can be written as $KB_{\pm}$ using 3.15, and hence $K$ acts chamber transitively on the twin building $\Delta_{\pm}$ associated to $G$. In particular
@@
This is a finite union of compact sets, see [HKM13, Corollary 3.10]. Recall also that the compact group $M$ is the maximal compact subgroup of the Lie group $T$ (cf. Proposition 3.7) and that $K_{l}^{\pm}\cap B_{\pm}\subseteq M$. Thus
@@
is a locally trivial fiber bundle by [Pal61] with compact fibers, and so are
@@
$k(g)\coloneqq\iota_{\pm}^{-1}\hskip-2.0pt\left(gAU_{\pm}\right)\;\text{ and }\;h(g)\coloneqq k(g)^{-1}g,$
where the maps $\iota_{\pm}$ are the same as in 3.18. Therefore we define the map
@@ 4.1 A criterion for symmetric spaces
Further we call a twin apartment $\Theta$-stable, if the twin building associated to $G$ is invariant as a set under the action of $\Theta$. Recall that $\Delta_{\pm}=G/B_{\pm}$, then we consider the action of $\Theta$ on the building as follows
@@ 4.2 Kac–Moody Symmetric Space
    and it can be deduced that the entries of $D$ are $\pm 1$. But by the positive semi-definite property only +1+1 is possible.
@@ 4.3 The Group Model
$\left\langle\tau\hskip-2.0pt\left(G^{\dagger}\right)\right\rangle=G^{\dagger}.$
@@
$=\ell_{g}\hskip-2.0pt\left(\Theta(g^{-1}h)K\right)$
$=\ell_{g}\hskip-2.0pt\left(s_{eK}\hskip-2.0pt\left(g^{-1}hK\right)\right)$
$=\left(\ell_{g}\circ s_{eK}\circ\ell_{g^{-1}}\right)\hskip-2.0pt(hK).$
@@
$\tau(G)^{2}C_{K}\hskip-2.0pt\left(G^{\dagger}A\right).$
@@ 4.4 Flats
$\mathfrak{a}\coloneqq\operatorname{span}_{\mathbb{R}}\hskip-2.0pt\left(\alpha^{\vee}_{1},\ldots,\alpha^{\vee}_{n},h_{n+1},\ldots,h_{2n-l}\right)\cong\mathbb{R}^{2n-l}.$
@@
The last step is to check that $gAK\subset G/K$ is closed. 3.19 tells that the multiplication $m\colon U_{\pm}\times A\times K\to G$ is a homeomorphism. This yields that the subspace $AK$ and $gAK$, for any $g\in G$, are closed in $G/K$ since the multiplication with an element induces a homeomorphism within the topological group. ∎
@@ 4.4.1 Flats in the Group Model
Recall that in the case $\mathbb{K}=\mathbb{C}$ $T$ is connected and in the case $\mathbb{K}=\mathbb{R}$ AA is the identity component of $T$. This indicates that $N_{K}(T)<N_{K}(A)$, based on the fact that the conjugation is continuous.
@@
To see that this map is well-defined, we first look at $N_{G}(T)$. We know by [FHHK20, Lemma 3.27] that $N_{G}(T)=A\ltimes N_{K}(T)$, since the proof of the quoted statement is given by 3.14 and the refined Birkhoff decomposition (see [Kum02, Theorem 5.2.3 (g)] or proof of 3.20). Together with 4.11 and [FHHK20, Lemma 3.27] follows that $N_{G}(T)$ stabilizes $AK$: Let $g\in N_{G}(T)$, then it can be written as $g=ak$ for $a\in A$ and $k\in N_{K}(A)$ and hence one calculates